6. The current Y across a resistor is a continuous uniform (-10, 10) a. Find P( IYI3) b. Find E[Y]. c. Find Var(Y).
Answer #1

Ans:

Given that Y is continuous uniformely distributed from -10 to 10.

a=-10

b=10

cdf:

P(Y<=y)=(y-a)/(b-a)=(y+10)/20

a)P(-3<Y<3)=P(Y<3)-P(Y<-3)=(3-(-3))/20=6/20=0.3

b)mean=E[Y]=(a+b)/2=0

c)Variance=Var[Y]=(b-a)^2/12=(10+10)^2/12=400/12=33.33